OpenStudy (razzanide):

Linear Systems Problem: The length of a rectangle is 9.7 cm more than 4 times the width. The perimeter of the rectangle is 91.4 cm. Write and solve a system of linear equations to represent the situation. I got my equations as l=9.7+4w and 2l+2w=91.4 (l=length, w=width) but now I'm not sure whether to use substitution or elimination?

OpenStudy (texaschic101):

I would use substitution. Just sub in 4W + 9.7 in for l in the other equation 2l + 2W = 91.4 2(4W + 9.7) + 2W = 91.4 can you finish from here ?

OpenStudy (razzanide):

2(4W + 9.7) + 2W = 91.4 8w+19.4+2w=91.4 10w=72 w=7.2 9.7+4(7.2)=l 9.7+28.8=l 38.5=l l=38.5; w=7.2

OpenStudy (texaschic101):

looks good...you can check it by subbing your info back into either equation 2l + 2W = 91.4 2(38.5) + 2(7.2) = 91.4 77 + 14.4 = 91.4 91.4 = 91.4 (correct)